User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair: Tiege

Attended: Katie, Simin, Mas, Patrick, Lucas

Apologies: Atifur, Jonny, Arwin, Ross


General Issues

  • How has Hobart 26m pressure/temps been holding up?
    • Keep an eye on this, be aware you may need to travel out to the telescope and pump the receiver.
    • 70K temp is the one to watch, along with pressure on log monitor (or run cryotemp a few times, take the lowest value).
  • Maser at Yg - VNC is no longer working!
    • Maybe need Brett to check it out.
  • Eremote control issues?
    • Make sure you check the troubleshooting section, might need to kill the ERCD process.
    • Can also use stream logs.
  • Module spreadsheet
    • Double check sum file and make sure the drive listed is suitable.
    • type 'sums_current' will open sum files for current experiment (also, 'sums_next')
  • Drifting clocks?
    • Keep an eye on this, notes on the wiki on how to fix it. reboot DBBC should fix as last resort.
  • LBA coming up next month.
    • Keep an eye on your emails.
  • be careful when Drudging from vex files, seems like it needs to be done on the relevant PC.

Handover Issues

* Disk usage preventing log files from being written (Yg)

  • Jamie has sorted this out hopefully.

* Drives stuck at Ke, this should be fixable on the drive control panel on timeke (or timeyg).

/home/www/auscope/opswiki/data/pages/operations/agenda2018_06_14.txt · Last modified: 2018/06/14 03:36 by Tiege McCarthy